# SpikeSift: A Computationally Efficient and Drift-Resilient Spike Sorting Algorithm

### v1

@@ Abstract
Spike sorting is a fundamental step in analyzing extracellular recordings, enabling the isolation of individual neuronal activity, yet it remains a challenging problem due to overlapping signals and recording instabilities, including electrode drift. While numerous algorithms have been developed to address these challenges, many struggle to balance accuracy and computational efficiency, limiting their applicability to large-scale datasets. In response, we introduce SpikeSift, a novel spike sorting algorithm designed to mitigate drift by partitioning recordings into short, relatively stationary segments, with spikes subsequently sorted within each. To preserve neuronal identity across segment boundaries, a computationally efficient alignment process merges clusters without relying on continuous trajectory estimation. In contrast to conventional methods that separate spike detection from clustering, SpikeSift integrates these processes within an iterative detect-and-subtract framework, enhancing clustering accuracy while maintaining computational efficiency. Evaluations on intracellularly validated datasets and biophysically realistic MEArec simulations confirm that SpikeSift maintains high sorting accuracy even in the presence of electrode drift, providing a scalable and computationally efficient solution for large-scale extracellular recordings.
@@ 1 Introduction
Neuronal communication occurs through action potentials—brief electrical impulses that encode and transmit information across neural circuits [1]. Analyzing these signals is essential for understanding brain function [2], diagnosing neurological disorders [3], and advancing neurotechnologies, including brain-computer interfaces [4]. While intracellular recordings provide highly precise measurements by inserting microelectrodes directly into neurons [5], their invasive nature limits their feasibility for large-scale and long-term studies [6]. Extracellular recordings offer a less invasive alternative by capturing voltage fluctuations in the surrounding medium. However, these signals reflect the combined activity of multiple nearby neurons [7], necessitating computational techniques to disentangle overlapping spike waveforms—a process known as spike sorting [8].
@@ 2.2.1 Template Formation
Once $c^{*}$ is designated, spike detection proceeds as oulined in Section 2.1.2. For each detected spike, a 2-ms waveform [8, 11] is extracted from the five electrodes closest to $c^{*}$, capturing the spike’s spatial distribution while maintaining computational efficiency [34], as shown in Figure 3a:
@@ 3.1 Paired Intracellular Recordings
The original dataset comprised 19 paired recordings; however, seven were excluded due to insufficient amplitude separation between intracellularly recorded spikes and background activity, which could have compromised the reliability of ground-truth validation. The remaining 12 recordings exhibited intracellular spikes with clear amplitude separation from background activity, ensuring a reliable reference for evaluating sorting accuracy. A summary of the recording durations and spike counts is presented in Table 1.
@@ 3.2 Biophysically Realistic Simulations
To thoroughly evaluate sorting performance, we generated recordings under four distinct conditions, each targeting a different aspect of SpikeSift’s robustness. A two-minute recording served as the benchmark, allowing direct comparison with existing spike sorting pipelines under standard experimental conditions. To assess performance on short recordings—where limited spike counts challenge clustering accuracy—we included a five-second dataset. For evaluating resilience to electrode motion, we generated two 10-minute recordings with different drift profiles: one with gradual drift at $5\mu$m/min [30] and another with discrete $25\mu$m jumps every two minutes to simulate abrupt electrode shifts. The jump magnitude was intentionally chosen to avoid alignment with inter-electrode spacing, preventing artificially easy correction. Together, these systematically varied scenarios provide a rigorous framework for assessing SpikeSift’s accuracy and stability in conditions that closely reflect real-world extracellular recordings.
@@ 4.2 Performance Evaluation
For simulated recordings, detected units were assigned to ground-truth neurons using the scoring metric from Kilosort, defined as:
@@ 4.3 Comparison with Other Algorithms
Mountainsort 5 segments recordings into discrete time blocks, performing clustering independently within each segment before merging clusters. It employs ISO-SPLIT [12], a clustering algorithm that recursively partitions data through unimodality tests, allowing it to adapt to complex waveform distributions without assuming predefined cluster shapes. To enhance drift adaptation, the block size was reduced from five minutes to one minute, allowing for finer adjustments. Initial trials showed that default PCA settings retained excessive dimensions, introducing noise rather than highlighting informative waveform differences. To improve feature extraction, the number of principal components per channel was reduced to 1, while the classifier nPCA was set to 20, enhancing cluster separability while minimizing unnecessary complexity.
@@ 4.4 SpikeSift Parameter Configuration
Figure 6: Sorting performance across different experimental conditions. The x-axis represents the number of correctly identified neurons, while the y-axis indicates either the number of spurious clusters or, in the second subfigure, the runtime in minutes. (a) Sorting results on a two-minute recording, comparing identified neurons and spurious clusters across methods. (b) Computational efficiency comparison, showing runtime as a function of the number of identified units. (c) Sorting accuracy in a short-duration recording (five seconds) (d) Sorting performance under gradual electrode drift ($5\mu$m/min). (e) Sorting accuracy in the presence of abrupt electrode jumps ($25\mu$m every two minutes).
@@ 5.1 Sorting Performance
Running on a single CPU core, SpikeSift consistently achieved a 10-fold speed advantage over Kilosort, Mountainsort, and SpyKING CIRCUS—despite their use of GPU acceleration or multi-threaded execution. For instance, Recording 2 from Table 1 was processed in just 13 seconds using SpikeSift, compared to 200 seconds for Kilosort and over 300 seconds for both Mountainsort and SpyKING CIRCUS. These performance differences persisted in simulated recordings, as shown in Figure 6b, where SpikeSift maintained a 20–40×40$\times$40 × speed advantage despite the additional complexity associated with the increased channel count.
@@
Table 2: Sorting performance across different values of $\kappa$.
@@ 5.2 Robustness to Electrode Drift
Sorting performance diverged most clearly in the five-second recording (Figure 6c), where limited spike counts posed a challenge for conventional methods. While accuracy declined sharply for other algorithms, SpikeSift maintained stable performance, demonstrating that its iterative detect-and-subtract framework remains effective even under sparse data conditions. This result highlights how SpikeSift overcomes the limitations that have traditionally hindered segmentation-based strategies in conventional pipelines.
This advantage becomes even more evident in the presence of electrode drift, as shown in Figures 6d and 6e. Mountainsort and SpyKING CIRCUS struggled to maintain consistent neuron identities over time due to limited or absent drift correction. SpyKING CIRCUS lacks any compensation mechanism, while Mountainsort processes data in fixed-duration blocks and applies a merging procedure that does not explicitly account for drift-induced shifts in waveforms.
@@
Table 3: Sorting performance across different values of $\lambda$.
@@
Table 4: Sorting performance under different drift magnitudes.
@@
5.3 Sensitivity Analysis
As shown in Table 2, lowering $\kappa$ to 7 enables SpikeSift to match Kilosort’s sensitivity, recovering 67 neurons, while producing only half as many spurious clusters (21 vs. 42). Conversely, increasing $\kappa$ to 13 eliminates all spurious detections while still recovering 50 valid neurons—demonstrating a level of specificity unmatched by competing methods.
These results underscore the flexibility of $L_{\min}$, the minimum segment duration. The default value of 10 seconds suffices to track fast drift, while the five-second benchmark in Figure 6c illustrates that shorter segments remain effective when required.
@@ Acknowledgments
This work was partially supported by the project “DeepTector”, funded by the the Special Account for Research Funds of the Aristotle University of Thessaloniki (project no. 10716).

### v2

@@ Abstract
Objective: Spike sorting is a fundamental step in analysing extracellular recordings, enabling the isolation of single‑neuron activity. However, it remains a challenging problem because extracellular traces mix overlapping spikes from neighbouring cells and are marred by recording instabilities such as electrode drift. Numerous algorithms have been proposed, yet many struggle to balance accuracy and computational efficiency, limiting their practicality for today’s large‑scale datasets. Approach: In response, we introduce SpikeSift, a spike‑sorting algorithm expressly designed to mitigate drift while running on standard CPUs. SpikeSift (i) partitions long recordings into shorter, relatively stationary segments, (ii) carries out spike detection and clustering simultaneously through an iterative detect‑and‑subtract scheme within each segment, and (iii) preserves neuronal identity across segments via a fast template‑alignment stage that dispenses with continuous trajectory estimation. Main results: Extensive validation on paired intracellularly validated datasets and on biophysically realistic MEArec simulations—covering elevated noise, diverse drift profiles, ultra‑short recordings and bursting activity—demonstrates that SpikeSift matches or exceeds the accuracy of state‑of‑the‑art methods while completing sorting an order of magnitude faster on a single desktop core. Significance: The combination of high fidelity, drift resilience, and modest computational demand renders SpikeSift broadly accessible while preserving data quality for downstream neurophysiological analysis.
@@ 1 Introduction
Neuronal communication occurs through action potentials—brief electrical impulses that encode and transmit information across neural circuits [1]. Analyzing these signals is essential for understanding brain function [2], diagnosing neurological disorders [3], and advancing neurotechnologies, including brain-computer interfaces [4]. While intracellular recordings provide highly precise measurements by inserting microelectrodes directly into neurons [5], their invasive nature limits their feasibility for large-scale and long-term studies [6]. Extracellular recordings offer a less invasive alternative by capturing voltage fluctuations in the surrounding medium; yet the resulting voltage traces represent the composite activity of nearby cells [7]. Work spanning more than two decades has shown that useful information can sometimes be extracted directly from this composite signal, without assigning spikes to individual neurons. Decoders that operate on threshold-crossing multi-unit activity or local-field potentials have reconstructed limb kinematics in motor cortex [8], and Bayesian models have decoded an animal’s spatial position from unsorted hippocampal activity [9]. Spectral–temporal embeddings of raw extracellular signals can likewise capture task-related dynamics without explicit clustering [10]. While such findings highlight the value of unsorted analyses, studies that probe synaptic connectivity, detailed population codes, or single-neuron biomarkers still demand cell-specific spike trains. In these cases the composite signal must be disambiguated through spike sorting, the computational attribution of each detected spike to its most probable source neuron [11].
@@ 2.1.1 Difference-of-Gaussians
The resulting band-passed trace is retained for both detection and clustering. Although band-pass filters can distort fine waveform shape and phase, such effects mainly hinder shape-based sorters [35, 36]. SpikeSift instead distinguishes units through the relative peak amplitudes recorded on neighbouring electrodes (see Section 2.2.2). Pilot checks showed that re-extracting snippets from the unfiltered signal increased noise in those amplitude ratios and produced no measurable gain in unit separability. Within this amplitude-centred framework, the filtered trace offers a practical balance between signal clarity and computational cost.
@@ 2.1.3 Adaptive Segmentation
Because the summed amplitude $S(t)$ is computed relative to the detection threshold, $H(t)$ captures changes in each unit’s excursion above baseline rather than absolute spike height. Moderate-amplitude neurons therefore influence boundary placement alongside larger ones, and the segmentation adapts automatically if the most prominent units diminish or drop out during a recording.
@@ 2.2.1 Template Formation
Once $c^{*}$ is designated, spike detection proceeds as outlined in Section 2.1.2. For each detected spike, a 2-ms waveform [11, 14] is extracted from the five electrodes closest to $c^{*}$, capturing the spike’s spatial distribution while maintaining computational efficiency [39], as shown in Figure 3a:
@@ 3.1 Paired Intracellular Recordings
The corpus originally comprised 19 paired recordings. Meaningful quantitative evaluation requires an intracellular trace whose spike events can be identified without ambiguity; otherwise, accuracy scores for any sorter would be unreliable. We therefore inspected the histogram of positive intracellular peaks in each file (16-bit integer samples). A recording was accepted when at least one unused integer value—an empty bin—separated the dense noise mode from a sparser band of spike amplitudes. Eleven of the twelve retained files showed a gap $>$$20\mu V$; the twelfth exhibited a shorter $2\mu V$ gap yet still displayed a clear valley. Gap size bore no relation to sorter accuracy—indeed, the file with the smallest gap was processed flawlessly by every method, whereas the first file that challenged all sorters contained a $170\mu V$ gap. The seven recordings that lacked any empty bin were excluded because noise and spike distributions overlapped continuously, making ground-truth labels unreliable. The twelve recordings that met the gap criterion constitute the benchmark summarised in Table 1.
@@ 3.2 Biophysically Realistic Simulations
To probe SpikeSift’s robustness under controlled yet challenging scenarios, 13 MEArec datasets were generated. The suite comprised a 2-minute baseline recording with $20\mu V$ RMS noise; two further 2-minute recordings with elevated noise levels of $30\mu V$ and $40\mu V$; a 5-second ultra-short trace to test performance with limited spike counts; three 10-minute recordings exhibiting distinct drift regimes—linear drift at $5\mu m\cdot min^{-1}$, a non-linear Gaussian-process trajectory spanning $50\mu m$ peak-to-peak, and abrupt displacements of 20–$25\mu m$ every two minutes; a set of five 1-minute recordings sweeping linear drift from 5 to $25\mu m\cdot min^{-1}$; and a 1-minute recording containing thirty intrinsically bursting neurons. Collectively, these datasets span key nuisance factors—noise amplitude, spike scarcity, drift geometry and magnitude, and non-stationary firing patterns.
@@ 4.2 Performance Evaluation
For simulated recordings, detected units were assigned to ground-truth neurons using the scoring metric from [17], defined as:
@@ 4.3 Comparison with Other Algorithms
Mountainsort 5 segments recordings into discrete time blocks, performing clustering independently within each segment before merging clusters. It employs ISO-SPLIT [15], a clustering algorithm that recursively partitions data through unimodality tests, allowing it to adapt to complex waveform distributions without assuming predefined cluster shapes. To enhance drift adaptation, the block size was reduced from five minutes to one minute, allowing for finer adjustments. Initial trials showed that default PCA settings retained excessive dimensions, introducing noise rather than highlighting informative waveform differences. To improve feature extraction, the number of principal components per channel was reduced to 1, while nPCA was set to 20, enhancing cluster separability while minimizing unnecessary complexity.
@@ 4.4 SpikeSift Parameter Configuration
Figure 6: Sorting performance under eight controlled conditions. Each bar shows correctly identified neurons (horizontal axis) against either spurious clusters (panels a, c–h; vertical axis) or runtime in minutes (panel b). (a) baseline 2-minute recording; (b) computational efficiency across methods; (c–d) elevated-noise recordings ($30\mu V$, $40\mu V$ RMS); (e) ultra-short 5-second recording; (f) slow linear drift ($5\mu m\cdot min^{-1}$); (g) non-linear Gaussian-process drift spanning $50\mu m$; (h) abrupt 20–$25\mu m$ jumps every 2 minutes.
@@ 5.1 Sorting Performance
Running on a single CPU core, SpikeSift consistently achieved a 10-fold speed advantage over Kilosort, Mountainsort, and SpyKING CIRCUS—despite their use of GPU acceleration or multi-threaded execution. For instance, Recording 2 from Table 1 was processed in just 13 seconds using SpikeSift, compared to 200 seconds for Kilosort and over 300 seconds for both Mountainsort and SpyKING CIRCUS. These performance differences persisted in simulated recordings, as shown in Figure 6b, where SpikeSift maintained a 20–40×40$\times$40 × speed advantage despite the additional complexity associated with the increased channel count. Across all datasets SpikeSift used less than 3 GB of RAM and wrote no temporary files, whereas competing pipelines created full on-disk duplicates of the filtered data and, for the 10-minute simulations, generated an additional $\approx$ 18 GB of intermediates. SpikeSift’s resource footprint is therefore essentially flat with recording length, while the alternatives scale at least linearly.
@@
Raising background noise from $20\mu V$ to $30\mu V$ and $40\mu V$ reduced unit yield for every sorter, yet their rank order did not change. At $40\mu V$ SpikeSift still recovered 25 ground-truth neurons while producing a single false cluster. Kilosort returned 27 true units but 36 spurious clusters, and SpyKING CIRCUS and Mountainsort lagged further, at 16 / 5 and 7 / 26 true / false units, respectively (Figures 6c and 6d.).
Table 2: Effect of detection threshold $\kappa$ on SpikeSift performance (2-minute baseline recording).
@@
A one-minute MEArec recording containing thirty intrinsically bursting neurons challenged each pipeline with strongly patterned firing. With its default threshold ($\kappa=10$) SpikeSift isolated 24 genuine units and produced no false clusters; Kilosort, SpyKING CIRCUS and Mountainsort reported 27 / 6, 16 / 5 and 7 / 26 true / false units, respectively. Lowering $\kappa$ to 9 elevated SpikeSift’s recall to 28 while preserving its zero-false-cluster record, matching Kilosort’s sensitivity but at markedly higher precision—a direct extension of the sensitivity–specificity trade-off shown in Table 2.
@@ 5.2 Robustness to Electrode Drift
Sorting performance diverged most clearly in the five-second recording (Figure 6e), where limited spike counts posed a challenge for conventional methods. While accuracy declined sharply for other algorithms, SpikeSift maintained stable performance, demonstrating that its iterative detect-and-subtract framework remains effective even under sparse data conditions. This result highlights how SpikeSift overcomes the limitations that have traditionally hindered segmentation-based strategies in conventional pipelines.
This advantage becomes even more evident in the presence of electrode drift, as shown in Figures 6f, 6g and 6h. Mountainsort and SpyKING CIRCUS struggled to maintain consistent neuron identities over time due to limited or absent drift correction. SpyKING CIRCUS lacks any compensation mechanism, while Mountainsort processes data in fixed-duration blocks and applies a merging procedure that does not explicitly account for drift-induced shifts in waveforms.
@@
Table 3: Effect of merging threshold $\lambda$ on SpikeSift performance (2-minute baseline recording).
@@
5.3 Sensitivity Analysis
As shown in Table 2, lowering $\kappa$ to 7 enables SpikeSift to match Kilosort’s sensitivity, recovering 67 neurons, while producing only half as many spurious clusters (21 vs. 42). Conversely, increasing $\kappa$ to 13 eliminates all spurious detections while still recovering 50 valid neurons—demonstrating a level of specificity unmatched by competing methods. These low false-positive counts show that the default pairing of a conservative detection threshold and a strict inter-cluster similarity test biases SpikeSift toward well-isolated single-unit clusters rather than broader multi-unit aggregates.
These results underscore the flexibility of $L_{\min}$, the minimum segment duration. The default value of 10 seconds suffices to track fast drift, while the five-second benchmark in Figure 6e illustrates that shorter segments remain effective when required.
Three recording contingencies may nonetheless warrant deliberate departures from the default settings. Substantial, short-interval drift ($>$$25\mu m$) is accommodated by widening the drift-alignment radius $D_{\max}$ and shortening the segment length $L_{\min}$ so that each segment remains approximately stationary. Recordings dominated by rarely firing neurons benefit from longer segments and a lower cluster-size threshold ($L_{\min}\uparrow$,$N_{\min}\downarrow$); a modest increase in the detection threshold $\kappa$ then suppresses noise that would otherwise contaminate sparsely populated clusters. Conversely, when low-amplitude spikes are the limiting factor, lowering $\kappa$ improves sensitivity, while a compensatory rise in $N_{\min}$ curbs the formation of noise-driven clusters.
Table 4: SpikeSift performance under increasing drift (five 1-minute recordings, 5–$25\,\mu$m/min).
@@ 6 Discussion
Looking ahead, SpikeSift is distributed as a lightweight Python package designed for scripted, reproducible workflows [48]. A graphical user interface that displays segment boundaries, cluster assignments, and merge operations—while allowing manual overrides—is a high-priority direction for future development. In the meantime, users can inject domain expertise through the existing programmatic hooks: a sorted recording can be divided with split_into_segments, rerun on selected portions via perform_spike_sorting with modified parameters, and curated or reconciled with merge_recordings and map_clusters. These tools allow experts to inspect, refine, and combine clusters directly within a scripted environment, offering a transparent and reproducible means of incorporating manual oversight.
@@ Acknowledgments
This work was partially supported by the project “DeepTector”, funded by the Special Account for Research Funds of the Aristotle University of Thessaloniki (project no. 10716).